OpenStudy (anonymous):

y - (y * .07) > or = 75,000

OpenStudy (vijay):

u mean y - (0.7*y) > or = 75,000

OpenStudy (anonymous):

Yes. I know the answer, but need to show it to teach my daughter

OpenStudy (anonymous):

no...sorry....(.07*y)

OpenStudy (vijay):

0.3y >= 75000 y>=250,000

OpenStudy (anonymous):

crap! sometimes you just look to make things difficult!

OpenStudy (anonymous):

thanks for ur help

OpenStudy (anonymous):

So y - .07 y = 1*y - .07y = (1 – .07) y = .93 y. Then you have .93 y > or = 75000. You can divide both sides by .93 to solve for y. You get y > or = 75000/ .93 = whatever the calculator says. Note that you don’t need to switch the signs on the inequality when you divide by a positive number. But if it had been like -.93y > or = than 75000, then dividing by -.93, you’d get that y < or = 75000/(-.93).